find the no such that one seventh of it is 2 more than 8.4

Grade:8

3 Answers

Arun
5 years ago
Dear Satyam
 
Let the number is x
 
the
 
x /7 = 2 + 8.4 = 10.4
 
x = 7 * 10.4 = 72.8
 
Regards
Arun (askIITians forum expert)
Sai Soumya
5 years ago
dear student , 
let the required number be x
ATQ, x/7=2+8.4=10.4
cross multiply
=x=7*10.4
=x=7.28
 
so the required answer is 7.28
Sayantan Garai
5 years ago
 
Dear Satyam
 
Let the number is x
 
then
 
x /7 = 2 + 8.4 = 10.4 
 
x = 7 × 10.4 = 72.8
(Answer)
------------------------------------